OpenStudy (anonymous):

2x+3 divided by 4= x over 4

OpenStudy (anonymous):

2x+3 /4 = x/4 So that means 2x+3=x So 2x=x-3 2x-x=-3 X=-3

OpenStudy (anonymous):

thank you so mucn, can you help me with more?